what does 9+3x=27 what does x equal??
Answer by venugopalramana(3286)

9+3x=27 ..to solve bring all unknowns to one side and knowns to the other side of equal to ,simplify and get the answer
3x=27-9=18
x=18/3=6
hence x=6...for a check substitute 6 for x in the given equation and verify.
9+3*6=9+18=27=right hand side ...ok